# YSO Jets Driven by Magnetic Pressure Generated through Stellar Magnetosphere–Disk Interaction

### v1

@@ Abstract
The origin of jets in young stellar objects (YSOs) remains a subject of active investigation. We present a 3D non-ideal magnetohydrodynamic simulation of jet launching in YSOs, focusing on the interaction between the stellar magnetosphere and the circumstellar disk. At the beginning of the simulation, the magnetosphere partially opens, forming two oppositely directed magnetic field regions: one threading the star and the other threading the inner disk. The latter dominates over the original disk field at small radii and contributes to launching a disk wind. In our model, the jet is launched from the interface between these regions by toroidal magnetic pressure generated along “two-legged” field lines, anchored at a magnetically dominated stellar footpoint and a mass-dominated point on the disk surface. Outflows are driven along these lines via a “load-fire-reload” cycle: in the “load” stage, differential rotation between the stellar and disk footpoints generates a toroidal magnetic field; in the “fire” stage, vertical gradients in toroidal field strength drive the outflow and release magnetic energy; and in the “reload” stage, magnetic reconnection between oppositely directed field lines resets the configuration, enabling the cycle to repeat. This process occurs rapidly and asynchronously across azimuthal angles, producing a continuous, large-scale outflow. From an energetic perspective, Poynting flux transports the toroidal field from the vicinity of the star into the polar cavity, powering the jet. Comparison with a disk-only model shows that the rotating stellar magnetosphere promotes bipolar jet launching by shaping a magnetic field geometry favorable to symmetric outflows.
@@ 1 Introduction
Star formation is often viewed as a collapse process, where a dense core, typically several to tens of thousands of AUs in size, collapses to form a young stellar object (YSO) consisting of a central star, a circumstellar disk, and, ultimately, planets. However, the reverse process, outflow, is also a fundamental aspect of the formation and evolution of YSOs. Outflows play a crucial role in shaping the chemical and dust distribution within individual star-forming systems (Y. Tsukamoto et al., 2021; L. Cacciapuoti et al., 2024; A. Morbidelli et al., 2024) and influence the formation of nearby stars and planets by regulating the dynamics of molecular clouds (P. André et al., 2014; D. Guszejnov et al., 2021; C.-H. Hsieh et al., 2023).
@@
The most energetic component of the outflow is the jet, a highly collimated stream of gas traveling at speeds exceeding $10^{7}~{}\mathrm{cm~{}s}^{-1}$, often extending thousands of AU or more in length (L. Podio et al., 2015; T. Ray & J. Ferreira, 2021). Recent observations have provided greater insight into the structure of jets and their surrounding environments (e.g., C. Codella et al., 2014; A. de Valon et al., 2020; C.-F. Lee et al., 2021; E. T. Whelan et al., 2024; K. D. Assani et al., 2024; J. A. López-Vázquez et al., 2024; S.-Y. Hsu et al., 2024; V. Delabrosse et al., 2024). Accompanying the jet is a slower, more massive outflow known as the disk wind, which typically surrounds the jet with a wider opening angle (B. Tabone et al., 2017; I. Pascucci et al., 2023). Since both the jet and disk wind significantly impact both local and global star formation processes, understanding their driving mechanisms is crucial to understanding star formation. A key open question is whether the jet and disk wind share a common driving mechanism, particularly what powers the jet itself.
@@
It is becoming a consensus that the jets and outflows are most likely magnetically driven. One commonly used model is the class of magnetocentrifugal models (F. H. Shu et al., 1995; R. Krasnopolsky et al., 1999; G. V. Ustyugova et al., 1999; R. Krasnopolsky et al., 2003; J. M. Anderson et al., 2005, 2006; X.-N. Bai & J. M. Stone, 2013), where the centrifugal force launches outflows as a result of angular momentum transportation along a magnetic field line. While the magneto-centrifugal model and its variants, including the “disk wind” (R. D. Blandford & D. G. Payne, 1982; R. E. Pudritz & C. A. Norman, 1983) and “X-wind” (F. Shu et al., 1994; F. H. Shu et al., 2000; H. Shang et al., 2002; M. J. Cai et al., 2008) models, can effectively explain the excessive angular momentum carried by the jet, how the exact launching conditions can be met under observational and theoretical constraints remains unclear (S. J. Desch et al., 2010; P. Bjerkeli et al., 2016).
@@
Another class of models is the Poynting flux-driven jet model, where the outflow is launched by a predominantly toroidal magnetic field that transfers energy and angular momentum to the gas (K. Shibata & Y. Uchida, 1986; R. V. E. Lovelace et al., 2001; R. V. E. Lovelace & M. M. Romanova, 2009; X. Guan et al., 2014). A related model is the “magnetic-tower” model (D. Lynden-Bell, 1996, 2003; Y. Kato et al., 2004; M. Nakamura et al., 2007), where the loop-like toroidal magnetic fields stack on top of each other, forming a magnetic loop tower. This class of models is usually more mass-loaded than the magneto-centrifugal model, resulting in a more gradual acceleration. Although this class of models provides an alternative method of acceleration, the stability of the required magnetic condition and the necessary external pressure to collimate the jet pose challenges to this model (R. Moll et al., 2008; M. Huarte-Espinosa et al., 2012).
@@
Either launching mechanism relies on energy and angular momentum extraction from its base, which is typically believed to be rooted in the inner disk or the star. One possibility is that the jet is driven by the disk alone. Since the YSO disk is expected to be magnetized, particularly during the early phase of disk evolution (J. Wurster & Z.-Y. Li, 2018; W. H. T. Vlemmings et al., 2019; M. Galametz et al., 2020; B. Lankhaar et al., 2022; H.-W. Yen et al., 2024; Y. Tu et al., 2024; S. Ohashi et al., 2025), the magnetic field in the disk could channel the disk’s rotational energy and momentum to a large distance, forming a jet. Previous studies have investigated this possibility (e.g. F. Casse & R. Keppens, 2002; S. X. Kato et al., 2002; J. M. Anderson et al., 2006; P. Tzeferacos et al., 2009; C. Fendt et al., 2011; C. Fendt & S. Sheikhnezami, 2013; D. Stepanovs & C. Fendt, 2016; W. Béthune et al., 2017; Z. Zhu & J. M. Stone, 2018; J. P. Ramsey & D. A. Clarke, 2019; G. Mattia & C. Fendt, 2020a, b; B. Mishra et al., 2020; J. Jacquemin-Ide et al., 2021; T. Jannaud et al., 2023; D. Takaishi et al., 2024; Y. Tu et al., 2025b, a; S. Takasao et al., 2025), but the jet-launching problem remains. One possibility is a jet can be launched by the disk alone, the jet is typically weak and variable, sometimes prone to being unipolar. Thus, it is unclear how general this jet-launching mechanism is in explaining observed outflows.
@@
One key ingredient missing in these disk-only jet models is the star, which can be connected to the inner disk through its magnetosphere. Observations have shown that YSOs could have a strong magnetosphere, and previous studies and simulations have primarily focused on how the magnetosphere influences the disk accreting onto the star (A. Bardou & J. Heyvaerts, 1996; K. A. Miller & J. M. Stone, 1997; M. M. Romanova et al., 2009, 2012; D. Lai, 2014; S. Takasao et al., 2019; M. M. Romanova et al., 2021, 2025; Z. Zhu et al., 2024; Z. Zhu, 2025). Outflow launching from the magnetosphere-disk boundary has been examined primarily in 2D (e.g., A. P. Goodson & R. M. Winglee, 1999; M. M. Romanova et al., 2005; R. V. E. Lovelace et al., 2013; M. Čemeljić et al., 2013; R. V. E. Lovelace et al., 2014), and less in 3D (e.g., M. M. Romanova et al., 2009; S. Takasao et al., 2022; Z. Zhu, 2025), showing the possibility of jet launching. Yet, how the magnetosphere participates in jet launching and how it interacts with the disk field remains uncertain. This paper focuses on modeling these key processes in a self-consistent 3D non-ideal magneto-hydrodynamic global model.
@@
This study is an extension of Y. Tu et al. (2025b) and Y. Tu et al. (2025a), who studied outflow launching using the disk gas and disk magnetic field alone. Y. Tu et al. (2025b) performed a 2D axisymmetric model and identified a key ingredient in jet launching, the “avalanche accretion stream” along the disk surface, which actively provides the energy and momentum to the jet-launching gas. Y. Tu et al. (2025a) extended the models to 3D and identified a light-loaded magnetic-pressure-driven jet model in a disk-only case, whose jet is highly variable and asymmetric in both disk hemispheres. In this study, we follow the disk setup in Y. Tu et al. (2025a) and include the stellar magnetosphere to examine the effect of the magnetosphere on jet launching.
@@
This paper is organized as follows. Sec. 2 describes the simulation setup and model parameters. Sec. 3 gives an overview of the simulation result and identifies the key components of the modeled jet-launching system. The jet-launching mechanism is described in detail in Sec. 4, followed by a discussion of this mechanism in the context of the star formation literature in Sec. 5. We conclude in Sec. 6.
2 Method
@@
$\frac{\partial\textbf{{B}}}{\partial t}=\nabla\times(\textbf{{v}}\times\textbf{{B}})-\frac{4\pi}{c}\nabla\times[\eta_{O}(\textbf{{J}}\times\hat{\textbf{{B}}})],$ (3)
@@
We use the Athena++ code (J. M. Stone et al., 2020) to solve the governing equations in Cartesian coordinates with static mesh refinement (SMR). The initial condition setup is similar to Y. Tu et al. (2025a), with the addition of the stellar magnetosphere. We refer the readers to Y. Tu et al. (2025a) for a detailed description of the hydro, ionization, and non-ideal MHD setup, which are briefly summarized here.
We adopt a power law $\rho=\rho_{0}(r/r_{0})^{-1.35}$ for the density of the disk midplane with $\rho_{0}=10^{-9}\mathrm{\ g\ cm^{-3}}$ and $r_{0}=6.3\times 10^{-2}\ \mathrm{au}$ following Y. Tu et al. (2025b). To model the truncation of the disk due to the stellar magnetosphere, the midplane density decreases as
@@
where $r_{\mathrm{trun}}=0.1$ au and $p=-10$. The remainder of the density is prescribed by assuming the disk is in isothermal hydrostatic equilibrium. To minimize numerical problems and ensure runability, we implement a similar density floor as in Y. Tu et al. (2025b):
@@
where $\rho_{f}=6\times 10^{-12}$ g cm-3, $r_{f}=0.015$ au. $q_{f}=-5$ allows a rapid drop in the density floor towards a large radius where a density floor is unnecessary. To speed up the simulation during the later stages, we increase the density floor coefficient $\rho_{f}$ to $2.4\times 10^{-11}$ g cm-3 after $t=1.0$ yr. The thermal and non-ideal MHD setup is identical to Y. Tu et al. (2025a), where thermal ionization of alkali metals keeps the inner disk ($r\lesssim 0.13$ au, the active zone) and the envelope in the ideal MHD limit, and Ohmic dissipation dominates the outer disk ($r\gtrsim 0.13$ au, the dead zone).
@@
Following Z. Zhu et al. (2024) and Z. Zhu (2025), we incorporate a dense “star” at the center of the simulation to minimize numerical diffusion of the stellar magnetosphere. This approach leverages the Cartesian coordinate system, which avoids the presence of voids within the simulation domain. The stellar radius is $R_{\star}=0.015~{}\mathrm{au}$ (about $3.2~{}R_{\odot}$), with a central density of $1.4~{}\mathrm{g~{}cm^{-3}}$. The remainder of the stellar density is calculated assuming the star is in isothermal hydrostatic equilibrium. To avoid numerical problems due to too large a density gradient at the stellar surface, we separate the star from the remainder of the simulation domain by setting a “fixing radius” $r_{\mathrm{fix}}=0.018$ au, within which the hydro quantities are reset to their initial values at each time step, while the magnetic field remains untouched. The passive scalars are set to 0 within $r_{\mathrm{fix}}$ at every time step. The gas in $r_{\mathrm{fix}}$ is assumed to rotate as a solid body with a rotational period of $12~{}\mathrm{days}$ (corresponding to the Keplerian rotation period at $0.1~{}\mathrm{au}$).
@@
where $\textbf{{A}}_{\mathrm{disk}}$ is the same as the vector potential used in Y. Tu et al. (2025a). The stellar magnetosphere is modeled following Z. Zhu et al. (2024), using a vector potential
@@
where $r_{c}=\max(r,r_{s}/2)$. $\bar{\textbf{{m}}}$ is chosen to be in the $-\hat{z}$ direction and of such magnitude that the initial magnetic field strength at the surface of the star is about 2000 Gauss. This choice of $\bar{\textbf{{m}}}$ ensures that the magnetic field of the magnetosphere points upward on the disk midplane, aligned with the direction of the initial open disk field.
@@
The computation domain and mesh refinement are chosen to best resolve the disk and the outflow-launching polar region while keeping a reasonable computational cost. The simulation domain spans ±$\pm~{}10~{}\mathrm{au}$ in the $\hat{x}$ and $\hat{y}$ directions, and ±$\pm~{}20~{}\mathrm{au}$ in the $\hat{z}$ direction. The root grid has a resolution of $64\times 64\times 128$, ensuring all cells are cubic. Nine levels of refinements are applied in the simulation, with the finest level applied in the disk (±$\pm 0.04~{}\mathrm{au}$ in the $\hat{x}$ and $\hat{y}$ directions, and ±$\pm 0.02~{}\mathrm{au}$ in the $\hat{z}$ direction), and in the outflow region (±$\pm 0.02~{}\mathrm{au}$ in the $\hat{x}$ and $\hat{y}$ directions, and ±$\pm 0.04~{}\mathrm{au}$ in the $\hat{z}$ direction). The remainder of the grid is generated automatically by applying the condition that adjacent meshblocks can only differ by at most 1 level of refinement. To reduce numerical diffusion within the star but not add too much computational cost, we apply two additional levels of refinement inside the star: level 10 covers ±$\pm 0.01~{}\mathrm{AU}$, and level 11 resolves the central ±$\pm 0.005~{}\mathrm{au}$. This refinement strategy keeps the stellar magnetosphere stable throughout the simulation. As a result of the mesh refinement, the finest cell size is $(0.0006~{}\mathrm{au})^{3}$ in the active domain, and $(0.00015~{}\mathrm{au})^{3}$ inside the star.
3 Result

### v2

# Modeling YSO Jets in 3D II: Accretion-Fed, Star-Anchored Poynting Jets in the Low-Density Polar Cavity Powered by Disk-Magnetosphere Interaction

@@ Abstract
The origin of jets in young stellar objects (YSOs) remains a subject of active investigation. We present a 3D magnetohydrodynamic simulation of jet launching in YSOs, focusing on the interaction between the stellar magnetosphere and the accretion disk. In our model, a fast, low-density bipolar jet is powered by disk-magnetosphere interaction and launched through the polar cavity that is mass-loaded from the disk rather than the star. Specifically, outflows are driven by toroidal magnetic pressure generated along “two-legged” field lines, anchored at a magnetically dominated stellar footpoint and a mass-dominated point on the (magnetically elevated) disk surface via a cyclic “load–fire–reload” process: in the “load” stage, differential rotation between stellar and disk footpoints generates toroidal magnetic pressure; in the “fire” stage, vertical gradients in the toroidal field accelerate plasma and transport Poynting flux into the polar cavity; in the “reload” stage, magnetic reconnection allows the cycle to repeat, reforming “two-legged” field lines. These field lines are not required to be fully reset to a dipolar loop configuration; it is only required that the disk-end be shallowly embedded in the (elevated) disk surface. This rapid, asynchronous process produces a continuous, large-scale outflow. The resulting magnetically dominated (Poynting) jet, accelerated by magnetic pressure within the low-density polar cavity, is distinct from the denser, slower disk wind launched through the classic magnetic-tower mechanism. Comparison with a disk-only model shows that the rotating stellar magnetosphere promotes bipolar jet launching by shaping a magnetic geometry favorable to symmetric outflows.
@@ I Introduction
Star formation is often viewed as a collapse process, where a dense core, typically several to tens of thousands of AUs in size, collapses to form a young stellar object (YSO) consisting of a central star, a circumstellar disk, and, ultimately, planets. However, the reverse process, outflow, is also a fundamental aspect of the formation and evolution of YSOs. Outflows play a crucial role in shaping the chemical and dust distribution within individual star-forming systems (Tsukamoto et al., 2021; Cacciapuoti et al., 2024; Morbidelli et al., 2024) and influence the formation of nearby stars and planets by regulating the dynamics of molecular clouds (André et al., 2014; Guszejnov et al., 2021; Hsieh et al., 2023).
@@
The most energetic component of the outflow is the jet, a highly collimated stream of gas traveling at speeds exceeding $10^{7}~\mathrm{cm~s}^{-1}$, often extending thousands of AU or more in length (Podio et al., 2015; Ray and Ferreira, 2021). Recent observations have provided greater insight into the structure of jets and their surrounding environments (e.g., Codella et al., 2014; de Valon et al., 2020; Lee et al., 2021; Whelan et al., 2024; Assani et al., 2024; López-Vázquez et al., 2024; Hsu et al., 2024; Delabrosse et al., 2024). Accompanying the jet is a slower, more massive outflow known as the disk wind, which typically surrounds the jet with a wider opening angle (Tabone et al., 2017; Pascucci et al., 2023). The difference in the properties of the jet and disk wind suggests these two outflows are launched from different locations and likely with different magnetic lever arms if they are driven magnetically (Lee et al., 2017; de Valon et al., 2020; Delabrosse et al., 2024). Since both the jet and disk wind significantly impact both local and global star formation processes, understanding their driving mechanisms is crucial to understanding star formation. A key open question is whether the jet and disk wind share a common driving mechanism, particularly what powers the jet itself.
@@
It is becoming a consensus that the jets and outflows are most likely magnetically driven. One commonly used model is the class of magnetocentrifugal models (Shu et al., 1995; Ferreira, 1997; Krasnopolsky et al., 1999; Ustyugova et al., 1999; Krasnopolsky et al., 2003; Anderson et al., 2005, 2006; Bai and Stone, 2013; Jacquemin-Ide et al., 2021), where the centrifugal force launches outflows as a result of angular momentum transportation along a magnetic field line. While the magneto-centrifugal model and its variants, including the “disk wind” (Blandford and Payne, 1982; Pudritz and Norman, 1983) and “X-wind” (Shu et al., 1994, 2000; Shang et al., 2002; Cai et al., 2008) models, can effectively explain the excessive angular momentum carried by the jet, how the exact launching conditions can be met under observational and theoretical constraints remains unclear (Desch et al., 2010; Bjerkeli et al., 2016).
@@
Another class of models is the Poynting flux-driven jet model, where the outflow is launched by a predominantly toroidal magnetic field that transfers energy and angular momentum to the gas (Shibata and Uchida, 1986; Lovelace et al., 2001; Lovelace and Romanova, 2009; Guan et al., 2014). A related model is the “magnetic-tower” model (Lynden-Bell, 1996, 2003; Kato et al., 2004; Nakamura et al., 2007), where the loop-like toroidal magnetic fields stack on top of each other, forming a magnetic loop tower. This class of models is usually more mass-loaded than the magneto-centrifugal model (Jacquemin-Ide et al., 2019), resulting in a more gradual acceleration. Although this class of models provides an alternative method of acceleration, the stability of the required magnetic condition and the external pressure necessary to collimate the jet pose challenges to this model (Moll et al., 2008; Huarte-Espinosa et al., 2012).
@@
We should note that, at the deepest level, the classic magneto-centrifugal and magnetic tower models are fundamentally the same, as discussed in, e.g., Ferreira (1997) and Spruit (2010, see his Sec. 3.2). Either launching mechanism relies on energy and angular momentum extraction from its base, which is typically believed to be rooted in the inner disk or the star. One possibility is that the jet is driven by the disk alone. Since the YSO disk is expected to be magnetized, particularly during the early phase of disk evolution (Wurster and Li, 2018; Vlemmings et al., 2019; Galametz et al., 2020; Lankhaar et al., 2022; Yen et al., 2024; Tu et al., 2024; Ohashi et al., 2025), the magnetic field in the disk could channel the disk’s rotational energy and momentum to a large distance, forming a jet. Previous studies have investigated this possibility (e.g. Casse and Keppens, 2002; Kato et al., 2002; Anderson et al., 2006; Zanni et al., 2007; Tzeferacos et al., 2009; Fendt et al., 2011; Murphy et al., 2010; Fendt and Sheikhnezami, 2013; Stepanovs and Fendt, 2016; Béthune et al., 2017; Zhu and Stone, 2018; Ramsey and Clarke, 2019; Mattia and Fendt, 2020a, b; Mishra et al., 2020; Jacquemin-Ide et al., 2021; Jannaud et al., 2023; Takaishi et al., 2024; Tu et al., 2025b, a; Takasao et al., 2025). However, when a well-ionized, relatively weakly magnetized inner disk is included self-consistently in our previous 3D global jet-formation simulations without a prescribed magnetic diffusivity, we found that the jet was relatively weak and often unipolar. It motivates us to include in the present study a key ingredient missing from our previous work: the stellar magnetosphere.
@@
Observations have shown that YSOs could have a strong magnetosphere, and previous studies and simulations have mainly focused on how the magnetosphere influences the disk accreting onto the star (Elsner and Lamb, 1977; Bardou and Heyvaerts, 1996; Miller and Stone, 1997; Bessolaz et al., 2008; Zanni and Ferreira, 2009; Romanova et al., 2009, 2012; Zanni and Ferreira, 2013; Lai, 2014; Takasao et al., 2019; Ireland et al., 2021; Romanova et al., 2021, 2025; Zhu et al., 2024; Zhu, 2025). Outflow launching from the magnetosphere-disk boundary has been examined primarily in 2D (e.g., Goodson and Winglee, 1999; Romanova et al., 2005; Lovelace et al., 2013; Čemeljić et al., 2013; Lovelace et al., 2014), and less in 3D (e.g., Romanova et al., 2009; Takasao et al., 2022; Zhu, 2025), showing the possibility of jet launching. Yet, how the magnetosphere participates in jet launching and how it interacts with the disk field remains uncertain. This paper focuses on modeling these key processes in a self-consistent 3D non-ideal magneto-hydrodynamic global model.
@@
This study extends the work of Tu et al. (2025b) and Tu et al. (2025a), who investigated outflow launching using the disk gas and the disk magnetic field alone. Tu et al. (2025b) performed a 2D axisymmetric model and identified a key ingredient in jet launching, the “avalanche accretion stream” along the disk surface, which actively provides the energy and momentum to the jet-launching gas. Tu et al. (2025a) extended the models to 3D and identified a light-loaded magnetic pressure-driven jet model in the disk-only case, whose jet is highly variable and asymmetric in both disk hemispheres [^1] We note that Zhu and Stone (2018) and Jacquemin-Ide et al. (2021) also performed disk-only simulations. Some similarities and differences between these works and ours are discussed in Tu et al. (2025b) (see their Section 4.1 in particular).. In this study, we follow the disk setup in Tu et al. (2025a) and include the stellar magnetosphere to examine the effect of the magnetosphere on jet launching.
@@
This paper is organized as follows. Sec. II describes the simulation setup and model parameters. Sec. III gives an overview of the simulation result and identifies the key components of the modeled jet-launching system. The jet-launching mechanism is described in detail in Sec. IV, followed by a discussion of this mechanism in the context of the literature in Sec. V. We conclude in Sec. VI.
II Method
@@
[…]
@@
We use the Athena++ code (Stone et al., 2020) to solve the governing equations in Cartesian coordinates with static mesh refinement (SMR). The initial condition setup is similar to Tu et al. (2025a), with the addition of the stellar magnetosphere. We refer the readers to Tu et al. (2025a) for a detailed description of the hydro, ionization, and non-ideal MHD setup, which are briefly summarized here.
We adopt a power law $\rho=\rho_{0}(r/r_{0})^{-1.35}$ for the density of the disk midplane with $\rho_{0}=10^{-9}\mathrm{\ g\ cm^{-3}}$ and $r_{0}=6.3\times 10^{-2}\ \mathrm{au}$ following Tu et al. (2025b). To model the truncation of the disk due to the stellar magnetosphere, the midplane density decreases as
@@
where $r_{\mathrm{trun}}=0.1$ au and p=10p=-10. The remainder of the density is prescribed by assuming the disk is in isothermal hydrostatic equilibrium. To minimize numerical problems and ensure runability, we implement a similar density floor as in Tu et al. (2025b):
@@
where $\rho_{f}=6\times 10^{-12}$ g cm-3, $r_{f}=0.015$ au, and $r$ is the spherical radius. $q_{f}=-5$ allows a rapid drop in the density floor towards a large radius where a density floor is unnecessary. To speed up the simulation during the later stages, we increase the density floor coefficient $\rho_{f}$ to $2.4\times 10^{-11}$ g cm-3 after $t=1.0$ yr. The thermal and non-ideal MHD setup is identical to Tu et al. (2025a), where thermal ionization of alkali metals keeps the inner disk ($r\lesssim 0.13$ au, the active zone) and the envelope in the ideal MHD limit, and Ohmic dissipation dominates the outer disk ($r\gtrsim 0.13$ au, the dead zone).
@@
Following Zhu et al. (2024) and Zhu (2025), we incorporate a dense “star” of $1~M_{\odot}$ at the center of the simulation to minimize numerical diffusion of the stellar magnetosphere. This approach leverages the Cartesian coordinate system, which avoids the presence of voids within the simulation domain. The stellar radius is $R_{\star}=0.015~\mathrm{au}$ (about $3.2~R_{\odot}$), with a central density of $1.4~\mathrm{g~cm^{-3}}$. The remainder of the stellar density is calculated assuming the star is in isothermal hydrostatic equilibrium. To avoid numerical problems due to too large a density gradient at the stellar surface, we separate the star from the remainder of the simulation domain by setting a “fixing radius” $r_{\mathrm{fix}}=0.018$ au, within which the hydro quantities are reset to their initial values at each time step, while the magnetic field remains untouched. The passive scalars are set to 0 within $r_{\mathrm{fix}}$ at every time step. The gas in $r_{\mathrm{fix}}$ is assumed to rotate as a solid body with a rotational period of $12~\mathrm{days}$ (corresponding to the Keplerian rotation period at $0.1~\mathrm{au}$).
@@
where $\textbf{{A}}_{\mathrm{disk}}$ is the same as the vector potential used in Tu et al. (2025a), where the disk is threaded by an open $\hat{z}-$direction magnetic field. The stellar magnetosphere is modeled following Zhu et al. (2024), using a vector potential
@@
where $r_{c}=\max(r,r_{s}/2)$. $\bar{\textbf{{m}}}$ is chosen to be in the $\hat{z}$ direction and of such magnitude that the initial magnetic field strength at the surface of the star is about 2000 Gauss. This choice of $\bar{\textbf{{m}}}$ ensures that the magnetic field of the magnetosphere points upward on the disk midplane, aligned with the direction of the initial open disk field.
@@
The computation domain and mesh refinement are chosen to best resolve the disk and the outflow-launching polar region while keeping a reasonable computational cost. The simulation domain spans $\pm~10~\mathrm{au}$ in the $\hat{x}$ and $\hat{y}$ directions, and $\pm~20~\mathrm{au}$ in the $\hat{z}$ direction. The root grid has a resolution of $64\times 64\times 128$, ensuring all cells are cubic. Nine levels of refinements are applied in the simulation, with the finest level applied in the disk ($\pm 0.04~\mathrm{au}$ in the $\hat{x}$ and $\hat{y}$ directions, and $\pm 0.02~\mathrm{au}$ in the $\hat{z}$ direction), and in the outflow region ($\pm 0.02~\mathrm{au}$ in the $\hat{x}$ and $\hat{y}$ directions, and $\pm 0.04~\mathrm{au}$ in the $\hat{z}$ direction). The remainder of the grid is generated automatically by applying the condition that adjacent meshblocks can only differ by at most 1 level of refinement. To reduce numerical diffusion within the star but not add too much computational cost, we apply two additional levels of refinement inside the star: level 10 covers $\pm 0.01~\mathrm{AU}$, and level 11 resolves the central $\pm 0.005~\mathrm{au}$. This refinement strategy keeps the stellar magnetosphere stable throughout the simulation. As a result of the mesh refinement, the finest cell size is $(0.0006~\mathrm{au})^{3}$ in the active domain, and $(0.00015~\mathrm{au})^{3}$ inside the star.
III Results
In this section, we present an overview of our simulation results and highlight key features of the jet structure. Because of the dynamic nature of the outflow launched in our model, we crudely divide the jet and disk wind using a simple velocity threshold: the jet is defined as gas with velocity >100>100 km s-1 away from the disk, and the disk wind as gas between $10$ and $100$ km s-1. The choice of $100$ km/s is somewhat arbitrary. Nevertheless, it separates well the low-density, fast-moving flow in the polar region from the much denser, more slowly moving surrounding flow in our particular simulation (see, e.g., Fig. 4a and Fig. 11, below). We have verified that the bulk of the jet material defined this way is super-fast-magnetosonic, especially at relatively large distances from the central star (see Fig.4a,b below), so that its launching should not be strongly modified by the jet interaction with its ambient medium.